OpenStudy (anonymous):

Which of the following points is a solution of the system shown? x + y ≥ 6 x ≥ 4 (-1, 7) (4, -2) (4, 2) i got c

OpenStudy (anonymous):

First, you turn the equations into the graphing form y=mx+b. The x+y+4 translates into y=-x+4 and the x-y=6 translates into y=x+6. Let's concentrate on the first equation.You graph this by first plotting b on your graph, which is the y-intercept, which would be positive 4. Then, you plot m , which is the slope of the line. The slope for this would be -1over positive 1. You start on the y intercept that you plotted and go down 1 number and then right one number.Repeat. You have plotted the first line. Now, for the second line, you first plot the y intercept again, which would be positive 6 this time. Next, you plot the slope which would be positive 1 over 1. To do this, you start at your y intercept and you go up one number and then to the right one number. Repeat. Don't forget to connect your dots for the graph. You have now successfully answered and graphed the equation!!!

OpenStudy (anonymous):

did that help?

OpenStudy (anonymous):

sorry one sec eating a taco

OpenStudy (anonymous):

am i correct than???????

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

thanks for the help